March BG Com Night 10, UT 2009 March 16

Observers: Dave Thompson (@LBT) NAgeorges, MJuette, WSeifert, ARakich in Tucson
Telescope Operator: Aaron Ceranski
SW Support: Paul Grenz
Telescope Support: TShih, Brian Curtis

Summary:

Dave got an excellent set of data for pointing model, and one data set for Jesper. Andrew got an excellent set of data for field aberration measurements. Seeing was excellent (best guide image ~ 0.27"). Lost 45 minutes for OSS problem. A couple of GCS restarts, lost several min each. GCS embedded SExtractor not yet ready for deployment. No telescope hardware problems. New pointing model apparently execellent. M3 home and return test showed no problems. Only complaint was that we ran out of night.

Details:

LBC Blue pumped and is cooling... T=212.1, P=7.24E-6 (Red T=147.6 P=2.63E-6).

UT 00:34 (45m before sunset): Sky basically clear with some thin cirrus. Tair=0.1C RH=42.3% Wind~4m/s. T(M1)=1.5C. Aaron reports the roof is clear and we can observe tonight!

The UPS on level one shows loads of: See also IT#404, the UPS may need to be load-balanced.

lbtmu05 seems very slow. Paul checked and said firefox was using a lot of memory. Stop and restart firefox and things seem to be OK now.

*Correct pointing and collimation*

UT01:47 Preset to BS9116, PARALLACTIC/TRACK to correct pointing, initial focus. IE changed from -31.0 to -15.04, CA does not need an offset.

UT01:57 preset ACTIVE on-axis, no star visible but the sky is still pretty bright (5000DN).
UT01:59 preset to off-axis guide star. Star not detected, but visible and out of focus. Applying focus global offset manually (+0.6 Z).
UT02:04 preset again, ACTIVE with on-axis guide star...successful. Collimating: 1311.71, 937.15, 325.36, 234.91, 411.78, ...347.00

02:12 started pointing log: /home/telescope/20090316.ptlog

Pointing model observations (TRACK mode) LUCIFER images=luci_20090315_NNNN. First 10 stars were deviating from the old rotator center by up to 30 arcsec, so we are taking 24 points now to make a new pointing model. The telescope is a few degrees warm, so if problems show up in this model we can take another set at the end of the night
UT Star LUCI
UT02:14 BS9116 0001
UT02:20 ACT0359 0002
UT02:26 WT10_406 0003
UT02:31 WT10_400 0004
UT02:35 ACT0344 0005
UT02:39 ACT0278 0006
UT02:43 WT10_176 0007
UT02:50 ACT0250 0008
UT02:54 ACT0095 0009
UT02:58 ACT0165 0010
UT03:05 ACT0271 0011
UT03:10 WT10_205 0012
UT03:14 WT10_324 0013
UT03:19 ACT0292 0014
UT03:23 ACT0337 0015
UT03:27 ACT0258 0016
UT03:32 WT10_218 0017
UT03:37 WT10_069 0018
UT03:42 WT10_112 0019
UT03:45 WT10_288 0020
UT03:51 WT10_303 0021

Hit X limit on M1, so we are stopping the pointing model data gathering.

Preset to BS9143, correct pointing: IE changed from -15.04 to -37.94, CA from 10.0 to 19.36. EL~48 and rising.

Rotator center:
UT03:51 PA=0 0023
UT03:51 PA=180 0024
UT03:51 PA=90 0025
UT03:51 PA=270 0026

GCS-AGW communication issue, WFS images failing to arrive:
Mar 15 21:28:06 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.

21:30:10.605 GCS stopped and restarted.

04:42 started taking an AGW transform calibration dataset. on BS9143 field

05:42 M3 timed out on the move...Paul looking into it.

23:07:35.760 Restarted GCS because it was failing to get WFS images from AZCam.
Mar 15 23:00:29 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Mar 15 23:02:00 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Mar 15 23:03:31 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Mar 15 23:05:05 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.
Mar 15 23:06:34 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 37716 ms.

Note on IIF GUI. The WFS RMS box is cleared and yellowed out on a preset, but when the WFS closes the loop it re-displays the previous valid measurement until the first WFS data at this position comes through. The box should stay yellowed out until the first valid rms at this position is calculated. Issue Trak this?

  • Take set of AGW aberration correction data

Field Aberration Measurements

1:18 Starting field aberration measurements.
1:32 Repository/Data/090316wfsc images 2-16 showing alot of bounce in the z4 - z6.
1:32 Switched GCS from using the internal centroiding to using sextractor.
1:34 Bounce seems to have dimished.
1:37 Sextractor seems to lose the star occasionally for no reason - maybe breaking up one star into many?

Persson's Star = BS9145
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
0 0.22 612.351 0 -188 2-25 , 29-30
6 1.15 477.9 343.83 -182 31-33
0 0.4 612.4 343.83 -183 34-35
18 0.95 506.19 too faint    
0 0.32 612.4 343.83 -184 36- 37
19 1.58 467.96 120.92 DOUBLE STAR!!!!!  
0 0.37 612.4 343.83 -185 38-40
7 0.83 479.5 194.32 -398 41-43
0 0.37 612.4 194.32 -399 44
6 1.15 477.9 343.83 -182 45-47
0 0.32 612.4 343.83 -184 48-49
7 0.83 479.5 194.32 -403 50-52
0 0.32 612.4 194.32 -405 53-59
swapping out of sextractor, back to Torsten algorithm

2:10 bouncing back and forward between gs 6 and 7, the only stars I can get in this field, to establish noise of measurment 2:11 sextractor failing on bright stars for no apparent reason. Needs more work before using with LUCIFER.

2:27 preset to 9152 (afer aborting a preset to 9143 that took us down to 40 el) and now have huge focus 4000nm wfsc60-61

2:45 big guiding jumps whith no apparent cause see 2:44:00 around then

3:15 getting 0.28" mean XY FWHM GS images (round too).

3:25 reset IE CA to -20 10

3:51 IDL failed to get WFSC image with GCS running in guide mode; restarting GCS seemed to fix the problem

4:05 GCS seemed to hang on preset, "waiting for handle block, never got to requesting full ccd readout"

4:10 got an RPC communication error from AGw. Restarting AGw. stopped and started AGw

4:30 determined that OSS had completely hung, all we did was open the OSS GUI. Paul restarted OSS and we could send presets again but failed at a different point. TPLC staus "heartbeat failed". See issuetrak 2069.

4:40 trying restart of MCSPU

4:50 that seemed to fix the heartbeat, and in hte process we had to re-home the tertiary, and the pointing was still good after that, so one test of homing the tertiary and going back to 25.9 degrees has been a success.

5:38 reset pointing IE CA -25 5

Persson's Star = BS9152
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
0 0.22 612.62 0 78 60-78
17 1.24 437.35 124.77 -306 84-86
0 0.1 612.37 124.77 -307 87-88
18 1.3 427.15 65.44 -250 89-91
0 0.26 612.43 65.44 -251 92-93
7 1.03 501.5 10.54 -198 94-96
0 0.26 612.43 10.54 -200 97-100
4 1.04 476.43 338.61 -530 101-103
0 0.26 612.43 338.61 -531 104-105
1 0.721 501.7 229.72 -425 106-108
0 0.26 612.43 229.72 -426 109-110
9 0.726 480.8 168.12 -366 111-113
0 0.26 612.43 168.12 -368 114-116
17 1.24 437.35 124.77 -326 117-119
0 0.26 612.43 124.77 -327 120-124
18 1.3 427.15 65.44 -273 125-127
0 0.26 612.43 65.44 -274 128-129
7 1.03 501.5 10.54 -220 130-132
0 0.26 612.43 10.54 -221 133-134
4 1.04 476.43 338.61 -191 135-137
0 0.26 612.43 338.61 -192 138-140
1 0.721 501.7 229.72 -445 141-143
0 0.26 612.43 229.72 -446 144-146: 147-149
9 .726 480.8 168.12 -389 150-152
0 0.26 612.43 168.12 -391 153-157

various events described above up to 4:40, now starting again

Guide Star x y PA Doug's Angle wfscimage000*
0 0.26 612.43 168.12 -408 158-164
17 1.24 437.35 124.77 -367 165-167
0 0.26 612.43 124.77 -368 168-170
18 1.3 427.15 65.44 -309 171-173
0 0.26 612.43 65.44 -310 174-176
7 1.03 501.5 10.54 -256 177-179
0 0.26 612.43 10.54 -257 180
4 1.04 476.43 338.61 -225 181-183
0 0.26 612.43 338.61 -226 184-185
1 .721 501.7 229.72 -478 186-188
0 0.26 612.43 229.72 -479 189: 190-194
9 .726 480.8 168.12 -421 196-198
0 0.26 612.43 168.12 -422 after repointing 199-202
17 1.24 437.35 124.77 -381 203-205
0 0.26 612.43 124.77 -381 206-207

5:50 going to BS9177 @ 66 deg EL, presetting around to find missing angles from previous data... need more at ~ 100 DA (GS 19 and up to 47 were good).

Persson's Star = BS9177
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
here comes the sun, da da da do da....

End time = 6:00

  • Repeat pointing measurements if earlier data show problems
  • M3 tests at end of night?

-- DavidThompson - 16 Mar 2009
Topic revision: r15 - 27 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback